Please complete all the required parts.
Part 1: CO / do Davis Company issued 10% annual rate, face value S600,000 bonds dated January 1, 20X3 on May 1, 20X3 to yield 8% annual market interest rate. The bonds pay interest every June 30 and December 31 and mature December 31 20X7. Assume the effective interest method. Required: A. Calculate the issue price of the bonds and set up an amortization schedule. B. Prepare the journal entries for the issuance and the first and second interest payments. C. Assume that after the second interest payment Davis bought back 40% of its bonds in the open market and retired them at a price of 98% of face value. What would be the journal entry for the buy back? Part 2: Assume the same information as in Part 1, except the bonds were issued on January 1, 20X3 and annual market interest rate is 12% . The bonds pay interest every June 30 and December 31 and mature December 31, 20X7. Assume the effective interest method. Required: alculate the issue price of the bonds and set up an amortization schedule. Prepare the journal entries for the issuance and the first, second, and third interest payments. Assume that after the third interest payment Davis bought back the bonds in the open market and retired them at a price of 102% of face value. What would be thejournal entry for the buy back? B. C.
Answer #1

As per policy, only one question is allowed to answer at a time, so answering Part 1 & its three parts here:

Part 1:
For half yearly interest, Period double & interest rate half
A) The issue of bond = half year coupon * PVIFA(Mrate,n) + Par value * PVIF(Mrate,n)
The issue of bond = (600000*10%*1/2) * PVIFA(4%,10) + 600000 * PVIF(4%,10)
The issue of bond =30000 * 8.111 + 600000 * 0.676 = $648930
AMORTIZATION SCHEDULE:
Interest paid Interest exp. Amortised amount Carrying Amount$
Jan 1,3 648930
Jun 30,3 30000 648930*4%=25957 4043 644887
Dec 31,3 30000 644887*4%=25795 4205 640683
Jun 30,4 30000 25627 4373 636310
Dec 31,4 30000 25452 4548 631762
Jun 30,5 30000 25270 4730 627033
Dec 31,5 30000 25081 4919 622114
Jun 30,6 30000 24885 5115 616999
Dec 31,6 30000 24680 5320 611679
Jun 30,7 30000 24467 5533 606146
Dec 31,7 30000 24246 5754 600392
B) Journal Entries:
Date Accounts Titles Debit $ Credit $
Jan 1,3 Cash 648930
Bond Payable 600000
Premium on Bond issue 48930
(being bond issue on premium)
Jun 30,3 Interest expense 25957
Premium on Bond issue 4043
Cash 30000
(being interest paid and premium adjusted)
Dec 31,3 Interest expense 25795
Premium on Bond issue 4205
Cash 30000
(being interest paid and premium adjusted)
C) Journal Entry of Buy Back of 40% of Bonds:
AMORTIZATION SCHEDULE:
Interest paid Interest exp. Amortised amount Carrying Amount$
Jan 1,3 259572
Jun 30,3 12000 10383 1617 257955
Dec 31,3 12000 10318 1682 256273
Journal Entry
Jan 1,4 Bond Payable 240000
Premium on Bond issue 16273 (256273-240000)
Cash 235200 (240000*98%)
Profit on redeemption of bonds 21073
(being 40% of bonds redeemed @ 98% and
premium adjusted )